The perimeter of a rectangle is 134m. The length is 7m more than twice the width. Find the Dimensions.
What is the length?
What is the Width?

Answer by mananth(16946)

let width be x
length = 2x+7
Perimeter = 2*(L+W)
2*(x+2x+7)=134
2*(3x+7)=134
6x+14=134
add -14
6x+14-14=134-14
6x=120
divide by 6
6x/6 = 120/6
x=20 the width.
length = 20*2+7= 47 m